For this equation: 2(x+3)=10 would x=5 and would it be 5+3+2=10?
Also if you check it that would be great!


-Ryan A.

Answer by stanbon(75887)

For this equation: 2(x+3)=10 would x=5?
------
2(x+3) = 10
Divide both sides by 2 to get:
x+3 = 5
Subtract 3 from both sides to get:
x = 2
===============
Cheers,
Stan H.
===============